# A Bayesian Optimization Approach to Machine Translation Reranking

### v1

@@ 3.2 Multi-fidelity BayesOpt
We also propose an extension to BayesOpt+GP for the setting where observations are available from a different but related proxy score function $s^{\prime}$. We refer to this as BayesOpt+GP+Proxy. $s^{\prime}$ is assumed to have non-trivial covariance with the scoring model $s$ and to be cheaper to evaluate. This is known as multi-fidelity BayesOpt in the literature, but while the multi-fidelity settings studied in (Kandasamy et al., 2016; Wu et al., 2020) use acquisition functions that may choose to evaluate lower-fidelity scores, we study a simpler setting: $m$ observations of $s^{\prime}$ are obtained at the start, and only $s$ may be evaluated during the BayesOpt loop. In the multi-fidelity setting, observations are made on $\langle y_{i},s_{i}\rangle$, a combination of a data point and scoring function, instead of the data point alone.
Our kernel for BayesOpt+GP+Proxy is the product of the RBF kernel from Section 3.1 and a kernel over score functions $f$:
$\mathcal{K}_{\mathrm{mult}}\big{(}\langle y_{i},f_{k}\rangle,\langle y_{j},f_{l}\rangle\big{)}=\\ \mathcal{K}_{\mathrm{MT}}(y_{i},y_{j})\mathcal{K}_{\mathrm{score}}(f_{k},f_{l}).$ (6)
$\mathcal{K}_{\mathrm{mult}}$ is a valid kernel because a product of two kernels defined on different spaces is also a kernel (Rasmussen and Williams, 2005). With $\mathcal{K}_{\mathrm{mult}}$, the covariance between two observations depends on both the difference between scoring functions and the distance between data points. This way, an observation influences the posterior for all other data points at all choices of scoring function, as long as the scoring functions are correlated. This formulation enables the use of any number of scoring functions, but in this work, we consider at most two: the main score and a proxy score.
We set $\mathcal{K}_{\mathrm{score}}(f_{k},f_{l})$ to be the empirical covariance between $f_{k}$ and $f_{l}$ measured over a validation set, where all scores are normalized per-instance so that in each instance, the scores of all candidates for a particular scorer have 0 mean and 1 variance. Then for each scoring function, concatenate all candidate scores across instances, and compare the resulting lists to obtain the covariance. Covariance is a valid kernel because the covariance calculation can be expressed as a dot product, and dot products are valid kernels.
@@ 3.3 Proxy scores
We train smaller scoring models to have high covariance with $s$ for use in BayesOpt+GP+Proxy. In this work, our scoring functions are based on the Comet referenceless quality estimation architecture (Rei et al., 2020), also known as CometKiwi. These models encode the source and hypothesis jointly with a bidirectional transformer. Activations from all transformer layers are pooled to form a fixed-size representation, which is passed to a feed-forward regression head. The vast majority of computation in this models is spent in the encoder. Thus, faster Comet models can be obtained by reducing the size of the encoder.
@@ 3.4 Candidate list generation
Figure 3: Performance of reranking methods measured as the average CometKiwi score of the selected candidate (left) and percentage of instances where the selected candidate had the highest score (right). The x-axis is the scoring budget. Legends show the normalized area under the curve of CometKiwi score of each method in brackets.
@@ 4 Experiments
We now discuss the details and findings of our Bayesian optimization experiments, followed by analysis of our trained proxy scoring models, concluding with runtime measurements. For exact values for figures in this section, see Table 3 in the Appendix. Appendix A contains extensive statistical significance tests.
For BayesOpt experiments, we grid search for the optimal value of RBF bandwidth parameter $w$ on the entire validation set, setting scoring budget $n=100$ and batch size $k=1$. While it is possible to optimize it for every unique combination of language pair, $n$, $k$, proxy scoring function, and $m$, we find that the results are not statistically significantly different within a range of settings. For simplicity, and to demonstrate the robustness of our methods, we use the same $w$ for all experiments.
We set $k{=}1$ in Sections 4.2 and 4.3 to demonstrate the effectiveness of BayesOpt+GP under ideal conditions, but since $k$ has a large impact on speed, we experiment with varying it in Section 4.5.
@@ 4.1 Models and datasets
For candidate generation, we use the 600M-parameter distilled NLLB model (Team et al., 2022) in all experiments. As the main scoring model, we use CometKiwi-22 (Rei et al., 2022b).
@@
For BayesOpt experiments, we randomly select 1000 and 500 source sentences per language pair from the WMT23 Metrics Shared Task dataset as validation and test set, respectively, for 7 language pairs: English-Czech, English-German, English-Japanese, English-Chinese, and the reverse directions of the latter 3 pairs.
@@ 4.2 BayesOpt+GP
The goal of reranking BayesOpt+GP is to improve the speed by only evaluating a subset of available candidates. We evaluate this through quality-cost tradeoff curves, where quality is determined by final selected candidate’s CometKiwi score, and cost is determined by the number of calls to the scoring function. As another measure of approximation quality, we also show the percentage of instances in which the actual best scoring candidate is returned. We devise several baselines to compare BayesOpt+GP against. Each is a strategy for selecting a subset of candidates to score from which the best scoring candidate is returned. The baselines are:
@@
UniqRandom simulates the effect of only sampling and scoring a smaller number of candidates while controlling the number of unique candidates. LogprobFirst{Avg,Sum} are included to verify whether more advanced methods indeed outperform simple subset selection using statistics obtained for free. HillClimbing is a heuristic iterative selection strategy which, like BayesOpt, is black-box and derivative-free (Conn et al., 2009).
@@ 4.3.1 Proxy score evaluation
LogprobsAvg 00.00s $\times$ $\times$ 0.191
LogprobsSum 00.00s $\times$ $\times$ -0.090
@@
Table 1: Benchmarking proxy models (Section 3.3) on speed and correlation with human judgments/CometKiwi using the WMT23 dataset. Speed is measured by runtime on a single A100-SXM4-40GB GPU per 10000 samples using maximum batch size. Correlation is measured with Kendall’s $\tau_{c}$ against human judgments and CometKiwi scores. CometKiwi correlation is taken over the provided targets in WMT23 (Test) and a synthetic dataset comprised of 200 samples per source sentence, deduped (Cands). Logprobs{Avg,Sum} is not evaluated on WMT23 targets because they are generated by other MT systems.
@@ 4.3.2 Reranking results
When $s^{\prime}$ is sufficiently fast and correlated with $s$, its knowledge can further improve the quality-cost tradeoff in BayesOpt+GP. Recall that BayesOpt+GP+Proxy initializes with $m$ evaluations of $s^{\prime}$. Figure 4 shows the quality-cost curve when all proxy scores are known. The relative performance when including proxy scores are ordered according to their correlation with CometKiwi as shown in Table 1; Distilled-M outperforms Distilled-S, and both outperform LogprobAvg. This demonstrates the importance of ensuring high correlation in the proxy score. The addition of LogprobAvg to BayesOpt+GP has little effect, showing that poorly correlated proxies are too noisy to help, and may even hinder performance. Beyond $n{=}70$, all methods achieve close to the maximum attainable score.
We also examine the effect of initializing with a fraction of proxy observations rather than all of them. For some choice of $m$, an appropriate baseline is to rank the top-$n$ candidates among the $m$ observed proxy scores. We call this ProxyFirst. The results when using Distilled-M and Distilled-S as proxies are shown in Figure 5. In both cases, the difference between BayesOpt+GP+Proxy and ProxyFirst is smaller when $m{=}200$ than when $m{=}50$, and this gap is smaller for Distilled-M. This is to be expected because as the covariance of $s$ and $s^{\prime}$ increases, using ProxyFirst with $m{=}200$ approaches standard full-list reranking. The marginal benefit of BayesOpt+GP+Proxy is more clear when $m{=}50$. In this case, proxy scores help to find promising candidates earlier, whereas ProxyFirst is a poor choice because it does not rank all of $\mathcal{C}$.
@@
Figure 4: Average CometKiwi score of the selected candidate (y-axis) for BayesOpt+GP+Proxy with different choices of proxy score.
Figure 5: Average CometKiwi score of the selected top candidate (y-axis) for BayesOpt+GP+Proxy with Distilled-S (left) and Distilled-M (right) compared to the ProxyFirst baseline. This figure disregards the additional compute costs for these proxy metrics in order to show the marginal score increase from proxy observations.
@@
Similarities $\times$ 001.24 001.24
BayesOpt+GP $\times$ 001.92 002.25
@@
Table 2: Runtimes for the full reranking baseline (AllComet), BayesOpt+GP, and BayesOpt+GP+S at settings where CometKiwi scores are roughly equal. Time given in seconds on a single A100-SXM4-40GB GPU for 350 samples.
4.4 Runtime
Our reranking algorithm significantly reduces actual runtime compared to scoring all candidates for a source sentence. We profile the full pipeline, from generating candidates to making a final selection, on three settings: (1) BayesOpt+GP with $n{=}90$, and (2) multi-fidelity BayesOpt+GP with 50 Distilled-S scores and $n{=}70$ (BayesOpt+GP+S), and 3) the baseline of evaluating CometKiwi on all candidates. $n,m$ are selected to balance the final scores of the two algorithms ($0.8213$ and $0.8211$ respectively, as shown in Table 3).
For the runtime calculations, we select 50 source sentences from each language pair and generate 200 candidates for each. For the baseline, we compute scores for all candidates with a batch size of 200. For our methods, we profile the additional steps required by BayesOpt: computing the kernel, computing the posteriors at each step, and optionally evaluating proxy scores. BayesOpt+GP(+S) uses batch size $k{=}10$, which does not affect scores compared to using $k{=}1$ (see Section 4.5). Memory bandwidth can be a major overhead in large neural networks, making it inefficient to run small batches. Since BayesOpt+GP obtains $k$ candidates per step, in order to use large batches, we process candidates for multiple instances in parallel.
Results are shown in Table 2. In all cases, candidate generation and CometKiwi calculations dominate the overall runtime. The extra cost from BayesOpt-related computations is compensated by the savings from reducing CometKiwi evaluations, despite similarity matrix computation being $\mathcal{O}(|\mathcal{C}|^{2})$, and matrix inversion for posterior calculation at each iteration being $\mathcal{O}(|\mathcal{C}|^{3})$. BayesOpt+GP+S reduces the runtime by further reducing the number of CometKiwi calculations to 70, with the cost of loading and running the Distilled-S proxy metric introducing minimal overhead.
@@ 6 Limitations
The optimization problem considered in this work is to maximize score from a scoring model. We show that BayesOpt is an effective optimizer, but we do not explore to what extent the optimization problem is flawed due to flaws in the scoring model. We refer to (Kocmi et al., 2024b) to understand what magnitude of score difference between systems is significant. However, the existence of “metric overfitting” when directly optimizing an evaluation metric is debated and may affect the interpretation of score differences (Fernandes et al., 2022; Wang et al., 2024b).
@@ Appendix A Statistical Significance
BayesOpt+GP with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
ProxyFirst 50 Distilled-S $\uparrow$ $\leftarrow$ $\uparrow$ $\uparrow$
ProxyFirst 50 Distilled-M $\uparrow$ $\leftarrow$ $\uparrow$ $\uparrow$
@@
BayesOpt+GP with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
ProxyFirst 50 Distilled-S $\uparrow$ $\leftarrow$ $\uparrow$ $\uparrow$
ProxyFirst 50 Distilled-M $\uparrow$ $\leftarrow$ $\uparrow$ $\uparrow$
BayesOpt+GP $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
ProxyFirst 50 Distilled-S $\leftarrow$ $\uparrow$ $\leftarrow$ $\uparrow$ $\uparrow$
ProxyFirst 50 Distilled-M $\leftarrow$ $\leftarrow$ $\uparrow$ $\uparrow$
@@
BayesOpt+GP with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$

### v2

@@ 3.2 Multi-fidelity BayesOpt
We also propose an extension to BayesOpt+GP for the setting where observations are available from a different but related proxy score function $s^{\prime}$. We refer to this as BayesOpt+GP+P. $s^{\prime}$ is assumed to have non-trivial covariance with the scoring model $s$ and to be cheaper to evaluate. This is known as multi-fidelity BayesOpt in the literature, but while the multi-fidelity settings of Kandasamy et al. (2016); Wu et al. (2020) use acquisition functions that may choose to evaluate lower-fidelity scores, we study a simpler setting: $\beta$ observations of $s^{\prime}$ are obtained at the start where $\beta>\alpha$, and only $s$ may be evaluated during the BayesOpt loop. In the multi-fidelity setting, observations are made on $\langle y_{i},s_{i}\rangle$, a combination of a data point and scoring function, instead of the data point alone.
Our kernel for BayesOpt+GP+P is the product of the RBF kernel from Section 3.1 and a kernel over score functions $f$:
$\mathcal{K}_{\mathrm{mult}}\big{(}\langle y_{i},s_{k}\rangle,\langle y_{j},s_{l}\rangle\big{)}=\\ \mathcal{K}_{\mathrm{MT}}(y_{i},y_{j})\mathcal{K}_{\mathrm{score}}(s_{k},s_{l}).$ (6)
$\mathcal{K}_{\mathrm{mult}}$ is a valid kernel because a product of two kernels defined on different spaces is also a kernel (Rasmussen and Williams, 2005). With $\mathcal{K}_{\mathrm{mult}}$, the covariance between two observations depends on both the difference between scoring functions and the distance between data points. This way, an observation influences the posterior for all other data points at all choices of scoring function, as long as the scoring functions are correlated. This formulation enables the use of any number of scoring functions, but in this work, we consider at most two: the main scorer $s$ and a proxy scorer $s^{\prime}$.
We set $\mathcal{K}_{\mathrm{score}}(s_{k},s_{l})$ to be the empirical covariance between $s_{k}$ and $s_{l}$ measured over a validation set, where all scores are normalized per-instance so that in each instance, the scores of all candidates for a particular scorer have 0 mean and 1 variance. Then for each scoring function, concatenate all candidate scores across instances, and compare the resulting lists to obtain the covariance. Covariance is a valid kernel because the covariance calculation can be expressed as a dot product, and dot products are valid kernels.
@@ 3.3 Proxy scores
We train smaller scoring models to have high covariance with $s$ for use in BayesOpt+GP+P. In this work, our scoring functions are based on the Comet referenceless quality estimation architecture (Rei et al., 2020), also known as CometKiwi. These models encode the source and hypothesis jointly with a bidirectional transformer. Activations from all transformer layers are pooled to form a fixed-size representation, which is passed to a feed-forward regression head. The vast majority of computation in this models is spent in the encoder. Thus, faster Comet models can be obtained by reducing the size of the encoder.
@@ 3.4 Candidate list generation
Figure 2: Left: Performance of reranking methods measured as the average CometKiwi score of the selected candidate. Beam search with beam size 5 achieves a score of 0.754 and is too low to be pictured here. Right: percentage of instances where the selected candidate had the highest score (right). The x-axis is the scoring budget. Legends show the normalized area under the curve of CometKiwi score of each method in brackets.
@@ 4 Experiments
We now discuss the details and findings of our Bayesian optimization experiments, followed by analysis of our trained proxy scoring models, concluding with runtime measurements. All run time values are measured on a A100-SXM4-40GB GPU. For exact values for figures in this section, see Table 3 in the Appendix. Appendix A contains extensive statistical significance tests.
For BayesOpt experiments, we grid search for the optimal value of RBF bandwidth parameter $w$ on the entire validation set, setting scoring budget $n=100$ and batch size $k=1$. While it is possible to optimize it for every unique combination of language pair, $n$, $k$, proxy scoring function, and $\beta$, we find that the results are not statistically significantly different within a range of settings. For simplicity, and to demonstrate the robustness of our methods, we use the same $w$ for all experiments.
In all experiments, we use $\alpha=10$ initial randomly scored candidates. We set $k=1$ in Sections 4.2 and 4.3 to demonstrate the effectiveness of BayesOpt+GP under ideal conditions, but since $k$ can have a large impact on speed, we experiment with varying it in Section 4.5.
@@ 4.1 Models and datasets
For candidate generation, we use the 600M-parameter distilled NLLB model (Team et al., 2022) in all experiments. For the main scoring model, we use CometKiwi-22 (Rei et al., 2022b).
@@
For BayesOpt experiments, we select the first 1000 and 500 source sentences per language pair from the WMT23 Metrics Shared Task dataset as the validation and test set, respectively, for 7 language pairs: English-Czech, English-German, English-Japanese, English-Chinese, and the reverse directions of the latter 3 pairs.
@@ 4.2 BayesOpt+GP
Figure 3: Average CometKiwi score of the selected top candidate (y-axis) for BayesOpt+GP+P with Distilled-S (left) and Distilled-M (right) compared to the ProxyFirst baseline. This figure disregards the additional compute costs for these proxy metrics in order to show the marginal score increase from proxy observations.
Figure 4: Average CometKiwi score of the selected candidate (y-axis) for BayesOpt+GP+P with different choices of proxy score.
The goal of reranking BayesOpt+GP is to improve the speed by only evaluating a subset of available candidates. We evaluate this through quality-cost tradeoff curves, where quality is determined by final selected candidate’s CometKiwi score, and cost is determined by the number of calls to the scoring function. As another measure of approximation quality, we also show the percentage of instances in which the actual best scoring candidate is returned. We devise several baselines with which to compare BayesOpt+GP. Each is a strategy for selecting a subset of candidates to score from which the best scoring candidate is returned. The baselines are:
@@
UniqRandom simulates the effect of iteratively sampling candidates until $n$ unique candidates are obtained. LogprobFirst{Avg,Sum} are included to verify whether more advanced methods indeed outperform simple subset selection using statistics obtained for free. HillClimbing is a heuristic iterative selection strategy which, like BayesOpt, is black-box and derivative-free (Conn et al., 2009).
@@ 4.3.1 Proxy score evaluation
LogprobsAvg 00.00s - - 0.191
LogprobsSum 00.00s - - -0.090
@@
Table 1: Benchmarking proxy models (Section 3.3) on speed and correlation with human judgments/CometKiwi using the WMT23 dataset. Speed is measured by runtime per 10000 samples using maximum batch size. Correlation is measured with Kendall’s $\tau_{c}$ against human judgments and CometKiwi scores. CometKiwi correlation is taken over the provided targets in WMT23 (Test) and a synthetic dataset comprised of 200 samples per source sentence, deduped (Cands). Logprobs{Avg,Sum} is not evaluated on WMT23 targets because they are generated by other MT systems.
@@ 4.3.2 Reranking results
When $s^{\prime}$ is sufficiently fast and correlated with $s$, it can further improve the quality-cost tradeoff in BayesOpt+GP. Recall that BayesOpt+GP+P initializes with $\beta$ evaluations of $s^{\prime}$. Figure 4 shows the quality-cost curve when all proxy scores are known, or $\beta=200$. The relative performance when including proxy scores correspond to their correlation with CometKiwi as shown in Table 1; Distilled-M outperforms Distilled-S, and both outperform LogprobAvg. This demonstrates the importance of ensuring high correlation in the proxy score. The addition of LogprobAvg to BayesOpt+GP has little effect, showing that poorly correlated proxies are too noisy to help and may even hinder performance. Beyond $n=70$, all methods achieve close to the maximum attainable score.
We also examine the effect of initializing with a fraction of proxy observations rather than all of them. For some choice of $\beta$, an appropriate baseline is to rank the top-$n$ candidates among the $\beta$ observed proxy scores. We call this ProxyFirst. The results when using Distilled-M and Distilled-S as proxies are shown in Figure 3. In both cases, the difference between BayesOpt+GP+P and ProxyFirst is smaller when $\beta=200$ than when $\beta=50$, and this gap is smaller for Distilled-M. This is to be expected because as the covariance of $s$ and $s^{\prime}$ increases, using ProxyFirst with $\beta=200$ approaches standard full-list reranking. The marginal benefit of BayesOpt+GP+P is more clear when $\beta=50$, where proxy scores help to find promising candidates earlier.
@@
4.4 Runtime
Our reranking algorithm significantly reduces actual runtime compared to scoring all candidates for a source sentence. We profile the full pipeline, from generating candidates to making a final selection, on three settings: (1) BayesOpt+GP with $n=90$, and (2) multi-fidelity BayesOpt+GP with 50 Distilled-S scores and $n=70$, and 3) the baseline of evaluating CometKiwi on all candidates. $n,\beta$ are selected to balance the final scores of the two algorithms ($0.8213$ and $0.8211$ respectively, as shown in Table 3).
For the runtime calculations, we select 50 source sentences from each language pair and generate 200 candidates for each. For the baseline, we compute scores for all candidates with a batch size of 200. For BayesOpt+GP methods, we profile the additional steps required: computing the kernel, computing the posteriors at each step, and evaluating proxy scores. BayesOpt+GP(+S) uses batch size $k=10$, which does not affect scores compared to using $k=1$ (see Section 4.5). Memory bandwidth can be a major overhead in large neural networks, making it inefficient to run small batches. Since BayesOpt+GP obtains $k$ candidates per step, in order to use large batches, we process candidates for multiple instances in parallel.
Results are shown in Table 2. In all cases, candidate generation and CometKiwi calculations dominate the overall runtime. The extra cost from BayesOpt-related computations is compensated by the savings from reducing CometKiwi evaluations, despite similarity matrix computation being $\mathcal{O}(|\mathcal{C}|^{2})$, and matrix inversion for posterior calculation at each iteration being $\mathcal{O}(|\mathcal{C}|^{3})$. BayesOpt+GP+P with Distilled-S reduces the runtime by further reducing the number of CometKiwi calculations to 70, with the cost of loading and running the Distilled-S proxy metric introducing minimal overhead.
@@ 4.4 Runtime
Similarities - 001.24 001.24
BayesOpt+GP - 001.92 002.25
@@
Table 2: Runtimes for the full reranking baseline (AllComet), BayesOpt+GP, and BayesOpt+GP+P with Distilled-S as proxy score at settings where CometKiwi scores are roughly equal. Time given in seconds per 350 instances.
@@ 6 Limitations
The optimization problem considered in this work is to maximize score from a scoring model. We show that BayesOpt is an effective optimizer, but we do not explore to what extent the optimization problem is flawed due to flaws in the scoring model. We refer to Kocmi et al. (2024b) to understand what magnitude of score difference between systems is significant. However, the existence of “metric overfitting” when directly optimizing an evaluation metric is debated and may affect the interpretation of score differences (Fernandes et al., 2022; Wang et al., 2024).
@@
Acknowledgements
Julius Cheng is supported by a scholarship from Huawei. Part of this work received support from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We thank the organizers of MT Marathon 2024, where the authors met and this work was conceived. We also thank Béni Egressy for useful discussions and Will Tebbutt for lending expertise on GPs.
@@ Appendix A Statistical Significance
BayesOpt+GP+P with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
BayesOpt+GP+P with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
BayesOpt+GP+P $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
BayesOpt+GP+P with LogprobAvg $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 200 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-S $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
BayesOpt+GP+P with 50 Distilled-M $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$